# Triple DES AES Advanced Encryption Standard

### v2

@@ 5.2. Microarchitectural Side-Channel/Covert-Channel Attacks
Dutta et al. (Dutta et al., 2021) develop two covert channels, enabling two malicious applications located on separate components (CPU and iGPU) to exchange confidential data through shared hardware resources. One channel exploits the shared Last-Level Cache (LLC) within Intel’s integrated GPU, while the other relies on the contention from the ring bus. Wang et al. (Wang et al., 2024a) develop cross-origin pixel stealing attacks in web browsers by exploiting the iGPU compression channel to induce data-dependent DRAM traffic and cache utilization. Given that the LLC is shared between the CPU and GPU, they demonstrate how to use LLC walk time to infer the GPU-induced CPU cache state changes. Wang et al. (Wang et al., 2021) extend the previous timing side-channel attack on AES implementation (Jiang et al., 2016) to the iGPU.
@@
Inferring sensitive information from GPU-based Neural Network (NN): Other attacks exploit side-channel leakage to infer information from ML models like NNs. Zhu et al. (Zhu et al., 2021) and Hu et al. (Hu et al., 2020) focus on the attack surface of the bus, particularly PCIe (as depicted in Figure 4(a)). Zhu et al. (Zhu et al., 2021) reverse-engineer the critical data structures (e.g., GPU command headers) offline and collect victim PCIe packets via a snooping device to reconstruct models in the online phase. The stolen DNN models have the same parameters and semantically identical architecture as the original ones. One potential mitigation for this attack is encrypting PCIe traffic. In contrast, Hu et al. (Hu et al., 2020) assume that the adversary cannot access data passing through buses, only the addresses, allowing attacks even with encrypted data. They conduct bus snooping attacks on discrete GPU’s memory and PCIe bus, obtaining kernel read/write access volume and memory address traces. Correlating extracted architectural events with model internal architectures. Another class of attacks occurs between two concurrent GPU applications, as shown in Figure 4(b). In (Naghibijouybari et al., 2017), they (i) implement website fingerprinting attack via the GPU memory utilization Application Programming Interface (API) or GPU performance counters; (ii) monitor user’s web activities during interactions or while typing characters on a keyboard through keystroke timing analysis; (iii) utilize a CUDA spy within the computational stack of GPU to deduce the NN structure by leveraging shared resources. In contrast to (Naghibijouybari et al., 2017), Wei et al. (Wei et al., 2020) consider fine-grained time-sliced sharing of applications on different VMs and exploit context-switching penalties on performance counters as leakage vector to recover the layer composition and hyperparameters of a DNN.
@@
Covert-channel attacks where both spy and victim concurrently run on a single GPU: The key point of a successful covert-channel attack is identifying the shared resources utilized by both the trojan and spy, often requiring extensive reverse engineering of GPU architectures. This process is necessary to uncover details about GPU components, such as the TLB and other shared resources, which can be exploited for covert communication. Naghibijouybari et al. (Naghibijouybari et al., 2017) first reverse-engineer the hardware block and warp-to-warp schedulers to establish the co-location, then learn contention on caches, functional units and memory, and finally construct covert channels based on these resources with an error-free bandwidth exceeding 4 Mbps. With this, they recover the number of neurons in an NN. Ahn et al. (Ahn et al., 2021b) initially reverse-engineer the structure of on-chip networks within modern GPUs and note that the hierarchical organization of the GPU leads to the sharing of interconnect bandwidth among adjacent cores. They then exploit the contention of the interconnect channel, leverage clock registers for sender-receiver synchronization and ultimately establish microarchitectural covert channels with a surprisingly high bandwidth of 24 Mbps. Nayak et al. (Nayak et al., 2021) reverse-engineer the details of the TLB hierarchy and the details of all TLB levels, such as the number of entries and associativity, of the NVIDIA 1080Ti. They create a covert timing channel via the shared L3 TLB using Prime+Probe. With MPS, they increase the bandwidth of the GPU’s L3 TLB-based covert channel to 81 Kbps. Unlike previous work relies on access to the native GPU APIs through CUDA or OpenGL for high-precision performance counters monitoring, Giner et al. (Giner et al., 2024) do not rely on WebGPU providing hardware timers. Instead, they use a shared memory buffer and a dedicated thread to increment a shared variable constantly, functioning as a timer. With this timer and their GPU-accelerated cache eviction set construction, they establish a Prime+Probe cache covert channel for the L2 cache with a bandwidth of 10.9 KB/s.
@@ 5.3. Physical Side-Channel Attacks
Zhan et al.(Zhan et al., 2022) identify the DVFS feature used in GPU as the underlying cause of an MMIO side-channel vulnerability. In scenarios where a victim employs an NVIDIA or AMD GPU without sharing, an attacker can monitor the victim’s activities and discern the visited webpages (i.e., a website fingerprinting attack). The key idea is capturing the fluctuating workload of a GPU during webpage rendering, which influences the MMIO signal it emits. These workload variations, coupled with the need to manage heat generation, fan noise, and minimize power consumption, lead to changes in the GPU’s performance levels (e.g., P-states in NVIDIA GPUs). Consequently, when a performance level switches on or off, distinct MMIO signals corresponding to the GPU’s memory clock frequencies appear or vanish in the spectrum. Remarkably, this attack can intercept MMIO signal leakage at distances of up to 6 meters, even through a wall. Liang et al. (Liang et al., 2022) conduct a similar far-field MMIO side-channel attack to infer the number of layers and their types, the number of kernels, sizes of layers and strides in an NN model.
@@ 5.4. Fault Injection Attacks
However, unlike encryption algorithms, where hardware fault injection attacks are highly effective, ML models, especially well-trained commercial DNN models, are inherently more fault-tolerant. This is because computation errors from random hardware faults are usually absorbed by later layers and do not greatly impact accuracy. Therefore, inducing hardware faults at specific timing points, optimizing the glitch voltage-frequency pair, and determining the appropriate fault duration are critical for generating controllable and impactful faults in ML models. To this end, Sun et al. (Sun et al., 2023) focus on searching sensitive targets of a CNN model, identifying proper parameters for timing and position, refining fault injection parameters, and ultimately reducing inference accuracy by an average of $69.1\%$. Santos et al. (Santos et al., 2021) propose a two-level fault injection framework combining the accuracy of Register Transfer Level (RTL) fault injection with the efficiency of software fault injection.
@@ 5.4.2. Countermeasures
6. Security of CPU TEEs
In this section, we introduce typical attacks against CPU TEEs. As there already exist survey papers focusing on reviewing existing attacks on CPU TEEs, and the aim of this survey is to explore potential attacks on GPU TEEs, we simply categorize the most common attacks based on their technical details and existing surveys (Fei et al., 2021; Nilsson et al., 2020; Cerdeira et al., 2020; Li, 2022) here.
6.1. Security of Intel SGX
6.1.1. Controlled Channel and Address Translation-based Attacks
Controlled channels attack, a concept introduced by Xu et al. (Xu et al., 2015) in 2015, exploits the privileges of the OS to monitor and manipulate page tables. This allows the attacker to uncover the memory access patterns of the enclave code, cf. also (Shinde et al., 2016). This attack is especially effective because, while Intel SGX enclaves protect sensitive computations from a compromised OS, the OS still manages memory and can observe page faults, thereby inferring details about the enclave’s operations. The SGX-Step (Van Bulck et al., 2017a) serves as a practical tool for, among others, executing such controlled channel attacks by configuring Advanced Programmable Interrupt Controller (APIC) timers, issuing interrupts, and tracking page-table entries. SGX-Step enables deterministic single-stepping of enclave execution, allowing for precise monitoring of memory access patterns over time.
6.1.2. Cache-based Attacks
6.1.3. Branch Prediction Attacks
6.1.4. Transient Execution Attacks
Transient execution attacks, which exploit out-of-order and speculative execution of instructions (summarized in section 2.2), have also been studied in the context of SGX. The aforementioned BTB and RSB attacks (Chen et al., 2019; Koruyeh et al., 2018) can also fall into this class of attacks, stemming from Spectre (Kocher et al., 2020). Additionally, the Foreshadow attack (Van Bulck et al., 2018), which targets SGX by employing Meltdown-like techniques, breaches enclave execution confidentiality. While MDS attacks (Canella et al., 2019; Van Schaik et al., 2019; Schwarz et al., 2019) also leverage speculative and out-of-order execution, they primarily exploit information leakage from various implementation-specific and undocumented intermediary buffers of the targeted microarchitecture. SGAxe (Van Schaik et al., 2020) and Crosstalk (Ragab et al., 2021) extend MDS to extract private keys from SGX enclaves.
6.1.5. Software-based Fault Injection and Side-Channel Attacks
6.1.6. Hardware-based Attacks
Apart from software-based fault injection attacks, Chen et al. (Chen et al., 2021; Chen and Oswald, 2023) explore hardware-level fault injection attacks on SGX enclaves. “VoltPillager” is a low-cost tool that injects messages on the Serial Voltage Identification (SVID) bus between the CPU and the voltage regulator, breaching the confidentiality and integrity of Intel SGX enclaves (Chen et al., 2021). Chen and Oswald later demonstrates that undervolting via the Power Management Bus (PMBus) can similarly break SGX enclave integrity, bypassing Intel’s countermeasures against previous software undervolting attacks (Murdock et al., 2020; Kenjar et al., 2020; Qiu et al., 2020). “MEMBUSTER” (Lee et al., 2020) introduces an off-chip attack that breaks SGX enclave confidentiality by snooping the memory bus (with high-spec, expensive lab equipment), extracting memory access patterns and increasing cache misses using techniques like critical page whitelisting and cache squeezing.
6.1.7. Interface-based Attacks
6.1.8. Countermeasures
(iii) Randomization: “SGX-Shield” (Seo et al., 2017) leverages ASLR to create a secure in-enclave loader that secretly randomizes memory space layout, hiding sensitive operations from attackers. Hosseinzadeh et al. (Hosseinzadeh et al., 2018) propose runtime control flow randomization, implemented as compiler extensions on llvm. “DR.SGX” (Brasser et al., 2019a) disrupts memory observations by permuting data locations.
(iv) Application/Source code design: Finally, developers can design enclave applications in a side-channel-resistant manner. Early works like “OBFUSCURO” (Ahmad et al., 2019) enforce operations with Oblivious RAM (ORAM) protection to prevent attackers from inferring access patterns. Several schemes (Ohrimenko et al., 2016; Poddar et al., 2020; Wang et al., 2022a) utilize oblivious primitives, implemented in x86_64 assembly, to design data-oblivious applications. These primitives operate exclusively on registers, which are private to the processor, ensuring that the contents hidden from code outside the enclave and any register-to-register data manipulation is data-oblivious.
6.2. Security of Arm TrustZone
6.2.1. Implementation Bugs
Cerdeira et al. (Cerdeira et al., 2020) examine bug reports obtained from public Common Vulnerabilities and Exposures (CVE) databases and vendor bulletin reports, categorizing them into validation bugs, functional bugs, and extrinsic bugs. These bugs are pervasive and often exploited to escalate privileges, enabling attackers to fully hijack kernels in devices from Qualcomm (Beniamini, 2016) and Huawei (Shen, 2015) or compromise client applications like Samsung Pay (Komaromy, 2018).
6.2.2. Cache-based Attacks
Given that the cache is shared between NW and SW in TrustZone-enabled processors, several studies capitalize on the contention arising from this cache coherency to extract sensitive information (Zhang et al., 2016c; Guanciale et al., 2016; Lipp et al., 2016; Zhang et al., 2016b; Cho et al., 2018). Several authors (Lipp et al., 2016; Zhang et al., 2016b; Cho et al., 2018) use the Prime+Probe method to monitor the cache activities within TrustZone.
6.2.3. Branch Prediction Attacks
6.2.4. Software-based Fault Injection Attacks
6.2.5. Countermeasures
In this section, we review and categorize countermeasures based on both previous survey (Cerdeira et al., 2020) and potential solutions from the summarized attack papers.
In addressing the architectural defences of Arm TrustZone, some research efforts (Sun et al., 2015; Brasser et al., 2019b) enhance isolation granularity between TEE components and reduce the amount of code running in SW, thereby lowering the risk of severe privilege escalation attacks within SW. Countermeasures (Jang et al., 2015; Jang and Kang, 2018) tackle the weaknesses of inadequate/weak authentication when accessing TEE resources from NW, and potentially insecure shared memory used for data exchange across the NW-SW boundary. Unlike Intel SGX, TrustZone lacks built-in on-chip memory encryption, prompting some works (Zhang et al., 2016a; Yun and Zhong, 2019) to design encryption strategies to bridge this gap. To further strengthen the verification of both TEE and TA binaries’ integrity and identity, several studies (Ferraiuolo et al., 2017; Lentz et al., 2018) introduced additional trusted computing primitives, such as enhanced remote attestation and sealed storage. To address implementation bugs, “RustZone” (Evenchick, 2018) extends TrustZone by implementing TAs using Rust, which inherently provides memory safety and thread safety.
Cache side-channel attacks can be mitigated by carefully implementing cryptographic algorithms in software (Guanciale et al., 2016) or using dedicated hardware (Lipp et al., 2016). Cache maintenance techniques can also be effective in mitigating leakage. For example, in terms of BTB attacks, flushing the shared microarchitectural structures when transferring between NW and SW can prevent attacks from targeting keys through the BTB. However, as observed in the Nexus 5X firmware by (Ryan, 2019), the L1 cache is not flushed, making Prime+Probe-based attacks still feasible.
6.3. Security of AMD SEV
6.3.1. Memory Encryption Issues
6.3.2. Software-based Fault Injection Attacks
Morbitzer et al. (Morbitzer et al., 2021) exploit the system’s lack of memory integrity protection to inject faults and execute arbitrary code within SEV VMs through I/O channels, using the hypervisor to locate and trigger the execution of encrypted payloads. However, SEV-SNP (Sev-Snp, 2020) can mitigate this issue through Reverse Map Table (RMP). Buhren et al. (Buhren et al., 2021) present a voltage fault injection attack enabling the execution of custom payloads on AMD secure processors, subsequently decrypting a VM’s memory by deploying custom SEV firmware. “CacheWarp” (Zhang et al., 2024) presents a new software-based fault attack targeting AMD SEV-ES and SEV-SNP, exploiting the possibility to revert modified cache lines of guest VMs to their previous (stale) state at the architectural level. This attack only requires interrupting the VM at a point chosen by the attacker to invalidate modified cache lines without writing them back to memory. More recently, Schlüter et al. (Schlüter et al., 2024a) introduce a novel attack where the hypervisor injects well-crafted, malicious #$\mathtt{\#VC}$ (a new exception by AMD facilitating communication between the VM and untrusted hypervisor) into a victim VM’s CPU, enabling the attacker to induce arbitrary behaviour in the VM. In another related attack, “HECKLER” (Schlüter et al., 2024b) shows that the hypervisor can inject malicious non-timer interrupts to compromise VMs. They leverage interrupt handlers with global effects, allowing manipulation of a VM’s register states to alter data and control flow. In general, SEV-SNP remains vulnerable to such software-based fault injection attacks.
6.3.3. Interrupt-driven Attacks
Similar to SGX-Step framework (Van Bulck et al., 2017a), SEV-Step (Wilke et al., 2024) introduces reliable single-stepping against SEV VMs. In addition to this functionality, SEV-Step facilitates access to common attack techniques, such as page fault tracking and cache attacks targeting SEV.
6.3.4. Countermeasures
7. Conclusions on future directions in GPU TEE
In this section, based on the relevant work discussed before, we sketch potential attacks that might be relevant to GPU TEEs in the future. We note that as GPU TEEs are not widely commercially available, it is not yet possible to practically evaluate these attack vectors, but believe that according vulnerabilities might be discovered in the future.

### v3

@@ 5.1. Architectural Attacks
Takeaway points. Architectural attacks often exploit leftover data in GPU memory due to missing or incomplete memory clearing. These attacks require only user-level access and affect both integrated and discrete GPUs, showing that such flaws are common across vendors. While many of these vulnerabilities have been patched in recent years, they highlight past weaknesses in GPU memory isolation.
@@ 5.2. Microarchitectural Side-Channel/Covert-Channel Attacks
Dutta et al. (Dutta et al., 2021) develop two covert channels, enabling two malicious applications located on separate components (CPU and iGPU) to exchange confidential data through shared hardware resources. One channel exploits the shared Last-Level Cache (LLC) within Intel’s integrated GPU, while the other relies on the contention from the ring bus. Wang et al. (Wang et al., 2024b) develop cross-origin pixel stealing attacks in web browsers by exploiting the iGPU compression channel to induce data-dependent DRAM traffic and cache utilization. Given that the LLC is shared between the CPU and GPU, they demonstrate how to use LLC walk time to infer the GPU-induced CPU cache state changes. Wang et al. (Wang et al., 2021) extend the previous timing side-channel attack on AES implementation (Jiang et al., 2016) to the iGPU.
@@
Takeaway points. Integrated GPUs introduce new security challenges due to their tight coupling with CPUs and shared microarchitectural resources, such as LLC and interconnects. This shared environment enables cross-component attacks, exploiting shared caches, compression channels, and GPU performance counters. These attacks are particularly concerning in web and mobile environments, where integrated GPUs are widely deployed and users have minimal control over low-level hardware access.
@@
Inferring sensitive information from GPU-based NN: Other attacks exploit side-channel leakage to infer information from ML models like NNs. Zhu et al. (Zhu et al., 2021) and Hu et al. (Hu et al., 2020) focus on the attack surface of the bus, particularly PCIe (as depicted in Figure 4(a)). Zhu et al. (Zhu et al., 2021) reverse-engineer the critical data structures (e.g., GPU command headers) offline and collect victim PCIe packets via a snooping device to reconstruct models in the online phase. The stolen DNN models have the same parameters and semantically identical architecture as the original ones. One potential mitigation for this attack is encrypting PCIe traffic. In contrast, Hu et al. (Hu et al., 2020) assume that the adversary cannot access data passing through buses, only the addresses, allowing attacks even with encrypted data. They conduct bus snooping attacks on discrete GPU’s memory and PCIe bus, obtaining kernel read/write access volume and memory address traces. Correlating extracted architectural events with model internal architectures. Another class of attacks occurs between two concurrent GPU applications, as shown in Figure 4(b). In (Naghibijouybari et al., 2017), they (i) implement website fingerprinting attack via the GPU memory utilization Application Programming Interface (API) or GPU performance counters; (ii) monitor user’s web activities during interactions or while typing characters on a keyboard through keystroke timing analysis; (iii) utilize a CUDA spy within the computational stack of GPU to deduce the NN structure by leveraging shared resources. In contrast to (Naghibijouybari et al., 2017), Wei et al. (Wei et al., 2020) consider fine-grained time-sliced sharing of applications on different VMs and exploit context-switching penalties on performance counters as leakage vector to recover the layer composition and hyperparameters of a DNN.
@@
Figure 5. Key steps of covert-channel attacks on a single GPU.
Covert-channel attacks where both spy and victim concurrently run on a single GPU: As summarized in Figure 5, the key point of a successful covert-channel attack is identifying the shared resources utilized by both the trojan and spy, often requiring extensive reverse engineering of GPU architectures. This process is necessary to uncover details about GPU components, such as the TLB and other shared resources, which can be exploited for covert communication. Naghibijouybari et al. (Naghibijouybari et al., 2017) first reverse-engineer the hardware block and warp-to-warp schedulers to establish the co-location, then learn contention on caches, functional units and memory, and finally construct covert channels based on these resources with an error-free bandwidth exceeding 4 Mbps. With this, they recover the number of neurons in an NN. Ahn et al. (Ahn et al., 2021b) initially reverse-engineer the structure of on-chip networks and find the hierarchical organization of the GPU leads to the sharing of interconnect bandwidth among adjacent cores. They then exploit the contention of the interconnect channel, leverage clock registers for sender-receiver synchronization and ultimately establish microarchitectural covert channels with a high bandwidth of 24 Mbps. Nayak et al. (Nayak et al., 2021) reverse-engineer the details of the TLB hierarchy and the details of all TLB levels, such as the number of entries and associativity, of the NVIDIA 1080Ti. They create a covert timing channel via the shared L3 TLB using Prime+Probe. With MPS, they increase the bandwidth of the GPU’s L3 TLB-based covert channel to 81 Kbps. Unlike previous work relies on access to the native GPU APIs through CUDA or OpenGL for high-precision performance counters monitoring, Giner et al. (Giner et al., 2024) do not rely on WebGPU providing hardware timers. Instead, they use a shared memory buffer and a dedicated thread to increment a shared variable constantly, functioning as a timer. With this timer and their GPU-accelerated cache eviction set construction, they establish a Prime+Probe cache covert channel for the L2 cache with a bandwidth of 10.9 KB/s.
@@
Takeaway points. ❶ Reverse engineering remains key to GPU attacks, highlighting both the limitations of security through obscurity and the lack of transparency from vendors that impedes systematic evaluation. ❷ Attack bandwidths have significantly improved and are now comparable to early CPU-based covert channels, reflecting the growing sophistication of GPU exploitation. ❸ Isolation mechanisms such as MIG may have flaws. Logical partitioning alone does not guarantee security and requires careful validation. ❹ Attacks leveraging NVLink and PCIe show that GPU interconnects are an emerging and under-protected attack surface, particularly in AI datacenter environments. ❺ ML workloads, especially DNNs, remain prime attack targets. Yet, the security of newer models such as LLMs and diffusion models across multi-GPU systems remains largely unexplored.
@@ 5.3. Physical Side-Channel Attacks
Zhan et al.(Zhan et al., 2022) identify the DVFS feature used in GPU as the underlying cause of an MMIO side-channel vulnerability. In scenarios where a victim employs an NVIDIA or AMD GPU, an attacker can monitor the victim’s activities and discern the visited webpages (i.e., a website fingerprinting attack). The key idea is capturing the fluctuating workload of a GPU during webpage rendering, which influences the MMIO signal it emits. These workload variations, coupled with the need to manage heat generation, fan noise, and minimize power consumption, lead to changes in the GPU’s performance levels (e.g., P-states in NVIDIA GPUs). Consequently, when a performance level switches on or off, distinct MMIO signals corresponding to the GPU’s memory clock frequencies appear or vanish in the spectrum. Remarkably, this attack can intercept MMIO signal leakage at distances of up to 6 meters, even through a wall. Liang et al. (Liang et al., 2022) conduct a similar far-field MMIO side-channel attack to infer the number of layers and their types, the number of kernels, sizes of layers and strides in an NN model.
Takeaway points.MMIO channels are a consistent and critical leakage source, acting as a bridge between CPU and GPU for both local and remote side-channel attacks. ❷ Access to internal sensors (e.g., temperature, power) varies by vendor. On platforms like Apple, where such access is unprivileged, this creates opportunities for exploitation. ❸ Remote physical attacks are becoming increasingly feasible, with some able to succeed from several meters away, even through walls, without privileged access. This broadens the threat model to include fully remote, passive attackers.
@@ 5.4. Fault Injection Attacks
However, unlike encryption algorithms, where hardware fault injection attacks are highly effective, ML models, especially well-trained commercial DNN models, are inherently more fault-tolerant. This is because computation errors from random hardware faults are usually absorbed by later layers and do not greatly impact accuracy. Therefore, inducing hardware faults at specific timing points, optimizing the glitch voltage-frequency pair, and determining the appropriate fault duration are critical for generating controllable and impactful faults in ML models. To this end, Sun et al. (Sun et al., 2023) focus on searching sensitive targets of a CNN model, identifying proper parameters for timing and position, refining fault injection parameters, and ultimately reducing inference accuracy by an average of $69.1\%$. dos Santos et al. (dos Santos et al., 2021) propose a two-level fault injection framework combining the accuracy of Register Transfer Level (RTL) fault injection with the efficiency of software fault injection.
@@
Takeaway points. ❶ Fault injection ‘can be triggered entirely through software by abusing out-of-spec voltage-frequency settings, exposing a lack of safeguards in GPU power management and its tight coupling with CPU controls. ❷ While DNNs are generally resilient to random faults, targeted injection at specific layers or operations can significantly reduce accuracy. Achieving this, however, requires precise control over timing, location, and fault parameters. ❸ The high memory bandwidth and parallelism of GPUs make them especially prone to Rowhammer attacks, as they can generate rapid memory access patterns that increase the risk of bit flips.
@@ 5.4.2. Countermeasures
6. Discussions on future directions in GPU TEE
In Section 4, we provided a taxonomy of existing GPU TEE designs, highlighting their architectural differences, security assumptions, and threat models. Building on this foundation, Sections 5 and the supplemental material (due to the limited space of this article) reviewed and compared a broad range of attacks targeting standalone GPUs and established CPU TEEs such as Intel SGX, Arm TrustZone, and AMD SEV. In this section, we extrapolate from these findings to identify potential attack vectors that may arise in the context of CPU TEEs, mapping them to the specific components and trust boundaries involved in these systems. We note that as GPU TEEs are not widely commercially available, it is not yet possible to practically evaluate these attack vectors, but we believe that according vulnerabilities might be discovered in the future.